OpenStudy (anonymous):

A mechanic uses a mechanical lift to raise a car. The car weighs 11,000 N. The lift raises the car 2.5 m. How much work is required to accomplish this task? A. 27,500 J B. 22,500 J C. 2.27 x 10-4 J D. 4400 J

OpenStudy (sidsiddhartha):

work done =force*distence=F*S here F=11000 N and S=2.5m WD=(11000*2.5)